OpenStudy (anonymous):

Sarah went to an electronics store, purchased 3 CDs and 5 DVDs, and spent $137. Michelle went to the same store.bought 2 CDs and 7 DVDs and spent $161. How Much does each CD sell for at the store?

OpenStudy (anonymous):

let C = cd and D = dvd. 3c + 5d = 137. Solve for one variable and plug into the other equation, 2c + 7 d = 161.

OpenStudy (anonymous):

let the cost of CD's be x and DVD's be y 3x+5y=137-------eq1 2x+7y=161-------eq2 now multiply 1st equation with 2nd and second equation with 3 6x+10y=274------1 6x+21y=483------2 now subtract equation 2 from 1 and u will get, -11y=-209 y=19 now substitute y in one of the equations 3x+5(19) = 137 3x = 137-95 x = 42/3 x = 14 so,the cost of a CD is 14 and DVD is 19